DETERMINE NO OF LIGHT EARS IN 1 METRE


 

Grade:10

3 Answers

Neeraj Sharma
12 years ago

149,598,500000m

Chetan Mandayam Nayakar
12 years ago

one light year=3*108*86400(number of seconds in a day)*365=9,460,800,000,000,000 metre,thus one metre=

(9,460,800,000,000,000)-1 =1.057*10^(-16)light year

bhanuveer danduboyina
12 years ago

Speed of light is 2.997 924 58 10^8 meters per second

multiply this with 86400 seconds per day

so you have 25'902'068'379'840 meters per day

multiply this with 365.24 days per year

Its equal to 9'460'471'000'000km per year

1 light year=9460800000000000 metres

PLEASE APPROVE MY ANSWER